Find an equation of the line passing trough the points (1,11) and (-2,2) using the point slope form: y-y1=m(x-x1)
Answer by richwmiller(17219)

plug in the values of y's and x's
11-2/(1-(-2))=m
9/3=3=m
11-y=3(1-x)
11-y=3-3x
8-y=-3x
8=y-3x
y-3x=8
y=3x+8